Tyler Company has the following information related to purchases and sales of one of its inventory items.

Date Description Units Purchased at Cost Units Sold at Retail
Sept. 1 Beginning inventory 400 units @ $12
Sept. 10 Purchase 600 units @ $13
Sept. 20 Sales 620 units @ $22
Sept. 25 Purchase 900 units at $15

Assume the company uses a perpetual inventory system.

Required:

Calculate ending inventory and cost of goods sold using the FIFO, LIFO, and average cost methods.
Answer #1

Calculate ending inventory and cost of goods sold

FIFO LIFO Average cost
Ending inventory (380*13+900*15) = 18440 (900*15+380*12) = 18060 26100-7812 = 18288
Cost of goods sold 26100-18440 = 7660 26100-18060 = 8040 620*12.60 = 7812